OpenStudy (anonymous):

pleaseee help! if you can at least help with one of them i would greatly appreciate it! factor completely, x^2+13x+36 factor completely, a^2-a-20 factor completely, x^2-5x-14 factor completely, b^2+12b+32

OpenStudy (anonymous):

x^2+13x+36 what two number MULTIPLIED gives you 36 but ADDED to give you 13?

OpenStudy (anonymous):

1) (x +9 )(x+4) because 4 *9 = 36 and 4+9 = 13

OpenStudy (anonymous):

do you know the next one?

OpenStudy (anonymous):

3) Factor: x^2 - 5x - 14? (x-7)(x+2)= 0 x=7, x=-2

OpenStudy (anonymous):

2) a^2 - a - 20 = a^2 - 5a + 4a - 20 = a (a-5) + 4(a-5) = (a+4)(a-5)

OpenStudy (anonymous):

(a -5 ) (a+4) because -5+4=-1 and -5 * 4 = -20

OpenStudy (anonymous):

4) (b^2)+12b+32=0 => (b^2)+8b+4b+32=0 => (b+8)(b+4)=0 => b= -8,-4.

OpenStudy (anonymous):

3) (x-7)(x+2) because -7 +2 = -5 and -7 *2 = -14

OpenStudy (anonymous):

thank you guys. i have another one- 8y^2+6y+1

OpenStudy (anonymous):

4) (b+4) (b+8) because 4+8 =12 and 4*8=32

OpenStudy (anonymous):

thanks sooo much sorry these are alot of problems.

OpenStudy (anonymous):

its fine

OpenStudy (anonymous):

no problem.....i love math

OpenStudy (anonymous):

haha that's good then i guess. can you do 8y^2+6y+1 and explain how to

OpenStudy (anonymous):

are we factoring it??

OpenStudy (anonymous):

I think what needs to happen is you get explained on how to do these problems, not just supplied answers. That is not the purpose of this site...

OpenStudy (anonymous):

i hope this is right... 8y^2 + 6y + 1 List the factors of a and c. a: 1 x 8, 2 x 4 c: 1 x 1 Since c is positive find factors from the list that will add up to the absolute value of b. 1, 1, 2, and 4 -----> (1 x 2) + (1 x 4) = 6 Arrange these four factors accordingly. Because both b and c are positive, both parentheses will have a plus sign. ( + )( + ) Ask yourself, "Which two factors with y as a variable will multiply to give me 8y^2?" Fill in the parentheses. (4y + )(2y + ) Use trial and error(by FOILing). (4y + 1)(2y + 1) Check answer. (4y)(2y) = 8y^2 (4y)(1) = 4y (2y)(1) = 2y (1)(1) = 1 Set up as a polynomial. 8y^2 + 4y + 2y + 1 Combine like-terms. 8y^2 + 6y + 1 So 8y^2 + 6y + 1 is equal to (4y + 1)(2y + 1).

OpenStudy (anonymous):

wow that was helpful lol. thank you

OpenStudy (anonymous):

lol your welcome...

OpenStudy (anonymous):

if you need help just ask ok

OpenStudy (anonymous):

okay! thanks

OpenStudy (anonymous):

@GamerGirl.1 15b^2-7b-2

OpenStudy (anonymous):

15b^2 - 7b - 2 (5b + 1) (3b - 2)

OpenStudy (anonymous):

thanks! 4a^2-8a-5

OpenStudy (anonymous):

4a^2 - 8a - 5 (2a - 5) (2a + 1)

OpenStudy (anonymous):

4y^2+25y+6

OpenStudy (anonymous):

4y^2 + 25y + 6 (4y + 1) (y + 6)

OpenStudy (anonymous):

omg thank you soo much! let me know if you ever need help with anything (except not math) lol

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

lol im good at math